Hello, I have a question on this problem:::
Could you please show me step by step how to get the answer. Thank You !
Simplify.
5-(8-3*4^3)
A)189
B)-179
C)1725
D)-195

Answer by Alan3354(69443)

Simplify.
5-(8-3*4^3)
= 5 - 8 + 3*4^3
= -3 + 3*64
= -3 + 192
= 189